# Time-Aware and Transition-Semantic Graph Neural Networks for Interpretable Predictive Business Process Monitoring  [^1] The full codebase is available at https://github.com/skyocean/TemporalAwareGNNs-NextEvent

### v1

@@ Abstract
Predictive Business Process Monitoring (PBPM) aims to forecast future events in ongoing cases based on historical event logs. While Graph Neural Networks (GNNs) are well suited to capture structural dependencies in process data, existing GNN-based PBPM models remain underdeveloped. Most rely either on short prefix subgraphs or global architectures that overlook temporal relevance and transition semantics. We propose a unified, interpretable GNN framework that advances the state of the art along three key axes. First, we compare prefix-based Graph Convolutional Networks(GCNs) and full trace Graph Attention Networks(GATs) to quantify the performance gap between localized and global modeling. Second, we introduce a novel time decay attention mechanism that constructs dynamic, prediction-centered windows, emphasizing temporally relevant history and suppressing noise. Third, we embed transition type semantics into edge features to enable fine grained reasoning over structurally ambiguous traces. Our architecture includes multilevel interpretability modules, offering diverse visualizations of attention behavior. Evaluated on five benchmarks, the proposed models achieve competitive Top-k accuracy and Damerau–Levenshtein scores without per-dataset tuning. By addressing architectural, temporal, and semantic gaps, this work presents a robust, generalizable, and explainable solution for next event prediction in PBPM.
@@ 1 Introduction
Predicting the next event in an ongoing business process is a cornerstone of operational intelligence across domains such as customer support, healthcare, finance, and manufacturing. Predictive Business Process Monitoring (PBPM) leverages historical event logs to anticipate future steps, enabling proactive decision making and real time optimization (Pasquadibisceglie et al., 2021; Zou et al., 2024). With the advancement of deep learning, graph neural networks have emerged as powerful tools for modeling complex event sequences (Harl et al., 2020). Despite increasing interest, existing graph neural network approaches for PBPM exhibit critical limitations in both their architectural design and their treatment of event and transition dynamics.
@@
First, most existing models rely either on short prefix based subgraphs (Dissegna and Di Francescomarino, 2025) or on global attention applied across entire traces. Prefix based approaches are computationally efficient but narrow in scope, often failing to capture long range dependencies. In contrast, global models provide richer context but tend to treat all events uniformly, without discounting distant and low signal history. Few studies compare these two paradigms or offer practical guidance on when one outperforms the other, leaving a methodological gap. Second, the integration of temporal information remains underexplored, particularly in graph neural network based models, which are well suited to encode both node level and edge level temporal signals. Temporal interactions can occur both locally—between consecutive events—and globally—between individual events and the overall sequence timeline. While local time differences have occasionally been used as static edge features in global models, their role is often limited and inconsistently applied. In contrast, global time alignment mechanisms, such as dynamic attention windows centered on the prediction point, are conceptually well motivated but remain almost entirely unexplored. To date, no study has systematically examined the comparative benefits of local versus global temporal integration within GNN architectures, despite their natural capacity to support both perspectives. Third, transitions between events, such as from complete request to close case, or from complete request to append close case, often carry meaningful semantic differences but are rarely modeled explicitly. This omission limits the model’s ability to differentiate between sequences that are structurally similar but behaviorally distinct. Fourth, many traditional machine learning benchmarks simplify event sequences by collapsing repeated events and ignoring substatus labels (e.g., lifecycle: transition). This reduction flattens the prediction space and overlooks common real world behaviors such as retries, escalations, and partial completions (Tama and Comuzzi, 2019). While recent deep learning models have begun to jointly predict both event type and substatus (Rama-Maneiro et al., 2023), this practice is still not widespread and remains absent from many comparative evaluations. Finally, GNNs are still infrequently applied to PBPM (Wang et al., 2025b) compared to recurrent or transformer based sequence models (Dissegna and Di Francescomarino, 2025). This underuse misses the opportunity to jointly model temporal progression and relational dependencies in event log data.
@@
To address these limitations, we propose a unified and interpretable graph neural network framework for predictive business process monitoring that advances the state of the art along three methodological axes:
@@
We validate our approach across five public benchmark datasets, where our models outperform existing baselines in both Top-k accuracy and sequence level alignment. Importantly, the interpretability analysis not only highlights predictive performance but also reveals the underlying decision logic, positioning the framework as both a forecasting solution and a diagnostic tool for process evaluation.
@@ 2 Related Work
Predictive Business Process Monitoring (PBPM) has progressed through successive methodological phases, each addressing prior limitations while introducing new challenges. The field has evolved from statistical models to neural architectures (Metzger et al., 2014; Park and Song, 2020) to better handle complex, long running processes. This progression contextualizes the emergence of graph neural network approaches and clarifies where significant gaps remain, particularly in temporal modeling and transition semantics.
@@ 2.1 From Prefix to Global Models
Early approaches in predictive business process monitoring modeled each case as a sequence prefix and applied deep learning models such as Long Short Term Memory networks (LSTM), Generative Adversarial Networks (GANs), Memory Augmented Neural Networks (MANN), and Convolutional Neural Networks (CNNs) (Tax et al., 2017; Camargo et al., 2019; Di Francescomarino et al., 2017; Taymouri et al., 2020; Khan et al., 2018; Pasquadibisceglie et al., 2019; Di Mauro et al., 2019; Sun et al., 2024) to predict upcoming events. While effective on short sequences, these models struggle with long traces containing loops or concurrency (Wang et al., 2025a), due to limited memory capacity and reliance on fixed prefix windows. To improve structural expressiveness, prefix based graph models such as Gated Graph Neural Networks (GGNNs) were proposed (Weinzierl, 2021), which encode direct event relationships within localized process segments. However, these models remain restricted to local context and are unable to dynamically reweight distant but informative events outside the prefix.
Recognizing these limitations, researchers have developed global sequence architectures that process entire traces simultaneously. Transformer based models (Bukhsh et al., 2021; Wuyts et al., 2024; Amiri Elyasi et al., 2024) leverage self attention mechanisms to capture long range dependencies, demonstrating improved performance over prefix based methods. Hierarchical variants (Ni et al., 2023) segment event sequences to capture multilevel semantics, addressing the complexity of nested or multigranular process structures. In parallel, global graph based approaches provide structured alternatives. For example, Pasquadibisceglie et al. (2024) apply heterogeneous Graph Attention Networks (GATs) to model interactions across multiple edge types, while Rama-Maneiro et al. (2023) combine recurrent neural networks with graph convolutional networks to jointly capture temporal and structural dependencies. Although these models improve performance, they still lack clearly defined and explainable mechanisms for the dynamics among node positions, temporal relevance and semantic transitions.
@@ 2.2 Temporal Dynamics in Process Prediction
While many models treat transitions between events uniformly, recent graph based approaches have begun to emphasize their semantic diversity and predictive importance. Pasquadibisceglie et al. (2024) incorporate multi type edges to distinguish between control flow and attribute based connections, demonstrating that transition semantics contribute significantly to prediction accuracy. Ayaz et al. (2024) apply graph autoencoders with edge conditioned convolutions to encode transition semantics, showing that different transition types require distinct representations. Lischka et al. (2025) preserve transition multiplicity using multi edge graphs and propose edge centric GNNs that treat transitions as first class entities. Despite these advances, the integration of semantic transition information into attention mechanisms remains rare and underexplored in current literature.
@@ 2.3 Interpretability in PBPM
As predictive models in predictive business process monitoring become more complex, interpretability has become essential for stakeholder trust and actionable insight. Early efforts relied on post hoc explainability methods such as surrogate models and feature attribution techniques (Mehdiyev and Fettke, 2021; El-Khawaga et al., 2022), but these typically offered limited visibility into the actual decision making processes of deep neural models. Recent research has shifted toward exploiting intrinsic interpretability within graph neural networks and attention based architectures. The PROPHET framework (Pasquadibisceglie et al., 2024) integrates GNNExplainer to identify influential nodes and edges, while Käppel et al. (2025) analyze attention weights in Transformer models to visualize control flow relevance across traces. Despite these developments, interpretability in GNN models for PBPM remains underdeveloped. Most existing work offers limited transparency into how temporal or structural signals shape predictions, and few approaches provide multilevel explanations that support both technical model validation and process level decision support.
@@ 4.1 Modeling Pipeline Overview
Our modeling framework evolves through a four stage pipeline, each stage progressively enhancing the models’ capacity to capture temporal and semantic dependencies in event sequences.
@@ 4.2.1 Graph Convolutional Networks
We adopt the GCNConv operator proposed by Kipf and Welling (2016), which uses a normalized adjacency matrix to stabilize message passing. The propagation rule at layer $l$ is adapted to our edge weighted, attributed graphs as follows:
@@ 4.3 Graph Attention Networks with Full Sequence Graphs
GATs enable flexible message passing by dynamically learning attention weights between connected nodes Velickovic et al. (2017). Unlike GCNs, which aggregate neighbor features via fixed edge weights, GATs learn attention scores that reflect the relative importance of each neighbor during aggregation.
@@ 4.3.4 Summary
The GAT models operate on full event sequence graphs, enabling the capture of long range dependencies across entire traces. Attention weights are conditioned on edge attributes, such as local time differences and optional transition type embeddings, allowing the model to emphasize contextually relevant events beyond immediate prefixes.
Variants incorporating dynamic temporal windows further apply exponential decay to attention scores based on each node’s temporal distance from the final event. This temporal weighting reflects the intuitive recency effect: recent events contribute more to prediction, while distant events have reduced influence. By jointly encoding what occurred (semantic transitions) and when it occurred (temporal decay), the model improves its ability to identify meaningful behavioral patterns in complex processes.
@@ 5.1 Data Description and Preprocessing
We evaluate our models on five benchmark event logs: BPI12, BPI12W, BPI13i, BPI13c, and Helpdesk. These datasets vary in temporal scale, process complexity, and attribute richness (Table 1), offering a comprehensive testbed for each stage of our modeling pipeline.
@@
Dataset #Seq Min(len) Max(len)222len = sequence length. t = time (minutes) between first and last events. #Attr(E, S) = number of event- and sequence-level attributes. med = median Med(len) Min(t) Max(t) Med(t) #Attr(E, S) #Evt #St #Cls33footnotemark: 3
@@
11footnotetext: When only one sequence exists at the maximum length, we report the second-longest length in brackets (representing the maximum length found in the test dataset).
A key observation from our dataset analysis is the discrepancy between the number of unique event types (#Evt) and the final classification labels (#Cls), which results from incorporating “lifecycle:transitions” as sub-status information (#St). For example, in the BPI13i and BPI13c datasets, the sub status categories (13 and 7, respectively) substantially increase the label space beyond the base event vocabulary (both with 4 activity types), leading to 3.25× and 1.75× larger output spaces. Given this expanded granularity in many real-world logs, we adopt the joint event and sub status prediction formulation, more commonly used in recent PBPM graph based models (Rama-Maneiro et al., 2021, 2023). This design better reflects the multi-dimensional nature of process states and supports more fine grained, actionable predictions.

### v2

@@ Abstract
Predictive Business Process Monitoring (PBPM) aims to forecast future events in ongoing cases based on historical event logs. While Graph Neural Networks (GNNs) are well suited to capture structural dependencies in process data, existing GNN-based PBPM models remain underdeveloped. Most rely either on short prefix subgraphs or global architectures that overlook temporal relevance and transition semantics. We propose a unified, interpretable GNN framework that advances the state of the art along three key axes. First, we compare prefix-based Graph Convolutional Networks(GCNs) and full trace Graph Attention Networks(GATs) to quantify the performance gap between localized and global modeling. Second, we introduce a novel time decay attention mechanism that constructs dynamic, prediction-centered windows, emphasizing temporally relevant history and suppressing noise. Third, we embed transition type semantics into edge features to enable fine grained reasoning over structurally ambiguous traces. Our architecture includes multilevel interpretability modules, offering diverse visualizations of attention behavior. Evaluated on six benchmarks with different decay factors, the proposed models achieve competitive Top-$k$ accuracy and Damerau–Levenshtein scores under a consistent global setup. A separate analysis of the decay coefficient $\lambda$ confirms its role in adapting attention, with effective values calibrated by temporal structure and sequence depth. By addressing architectural, temporal, and semantic gaps, this work presents a robust, generalizable, and explainable solution for next event prediction in PBPM.
@@ 1 Introduction
Predicting the next event in an ongoing business process is a cornerstone of operational intelligence across domains such as customer support, healthcare, finance, and manufacturing. Predictive Business Process Monitoring (PBPM) leverages historical event logs to anticipate future steps, enabling proactive decision making and real time optimization (pasquadibisceglie2021multi; zou2024prediction). With the advancement of deep learning, graph neural networks have emerged as powerful tools for modeling complex event sequences (harl2020explainable). Despite increasing interest, existing graph neural network approaches for PBPM exhibit critical limitations in both their architectural design and their treatment of event and transition dynamics.
@@
First, most existing models rely either on short prefix based subgraphs (dissegna2025graph) or on global attention applied across entire traces. Prefix based approaches are computationally efficient but narrow in scope, often failing to capture long range dependencies. In contrast, global models provide richer context but tend to treat all events uniformly, without discounting distant and low signal history. Few studies compare these two paradigms or offer practical guidance on when one outperforms the other, leaving a methodological gap. Second, the integration of temporal information remains underexplored, particularly in graph neural network based models, which are well suited to encode both node level and edge level temporal signals. Temporal interactions can occur both locally—between consecutive events—and globally—between individual events and the overall sequence timeline. While local time differences have occasionally been used as static edge features in global models, their role is often limited and inconsistently applied. In contrast, global time alignment mechanisms, such as dynamic attention windows centered on the prediction point, are conceptually well motivated but remain almost entirely unexplored. To date, no study has systematically examined the comparative benefits of local versus global temporal integration within GNN architectures, despite their natural capacity to support both perspectives. Third, transitions between events, such as from complete request to close case, or from complete request to append close case, often carry meaningful semantic differences but are rarely modeled explicitly. This omission limits the model’s ability to differentiate between sequences that are structurally similar but behaviorally distinct. Fourth, many traditional machine learning benchmarks simplify event sequences by collapsing repeated events and ignoring substatus labels (e.g., lifecycle: transition). This reduction flattens the prediction space and overlooks common real world behaviors such as retries, escalations, and partial completions (tama2019empirical). While recent deep learning models have begun to jointly predict both event type and substatus (rama2023embedding), this practice is still not widespread and remains absent from many comparative evaluations. Finally, GNNs are still infrequently applied to PBPM (wang2025hgcn) compared to recurrent or transformer based sequence models (dissegna2025graph). This underuse misses the opportunity to jointly model temporal progression and relational dependencies in event log data.
@@
To address these limitations, we propose a unified and interpretable graph neural network framework for PBPM that advances the state of the art along three methodological axes:
@@
We validate our approach across six public benchmark datasets, where our models outperform existing baselines in both Top-k accuracy and sequence level alignment. Importantly, the interpretability analysis not only highlights predictive performance but also reveals the underlying decision logic, positioning the framework as both a forecasting solution and a diagnostic tool for process evaluation.
@@ 2 Related Work
Predictive Business Process Monitoring (PBPM) has progressed through successive methodological phases, each addressing prior limitations while introducing new challenges. The field has evolved from statistical and classical machine learning models (di2022predictive) (e.g., decision trees (de2016general), random forests (di2016clustering), and regressions (van2012process)) to neural architectures (metzger2014comparing; park2020predicting) to better handle complex, long running processes. More recent studies highlight the rapid adoption of deep learning and graph based approaches (deng2024enhancing), reflecting a shift toward models that can capture temporal irregularities (hennig2025leveraging), structural dependencies, and contextual attributes at scale (fioretto2025comparative). This progression contextualizes the emergence of graph neural network approaches and clarifies where significant gaps remain, particularly in temporal modeling and transition semantics.
@@ 2.1 From Prefix to Global Models
Early approaches in PBPM modeled each case as a sequence prefix and applied deep learning models such as Long Short Term Memory networks (LSTM), Generative Adversarial Networks (GANs), Memory Augmented Neural Networks (MANN), and Convolutional Neural Networks (CNNs) (tax2017predictive; camargo2019learning; di2017eye; taymouri2020predictive; khan2018memory; pasquadibisceglie2019using; di2019activity; sun2024next) to predict upcoming events. While effective on short sequences, these models struggle with long traces containing loops or concurrency (wang2025comprehensive), due to limited memory capacity and reliance on fixed prefix windows. To improve structural expressiveness, prefix based graph models such as Gated Graph Neural Networks (GGNNs) were proposed (weinzierl2021exploring), which encode direct event relationships within localized process segments. However, these models remain restricted to local context and are unable to dynamically reweight distant but informative events outside the prefix.
Beyond accuracy, prefix-based models also faced issues of prefix bias, where early predictions benefited from dense supervision but later stages suffered from data sparsity (pauwels2021incremental). Several works attempted to mitigate this with prefix augmentation (pfeiffer2025learning) or sampling strategies (rama2024towards), yet performance remained sensitive to prefix length (klinkmuller2018towards) and trace variability (stevens2025generating). Moreover, while CNNs and GAN (hoffmann2021progan) based models captured local sequential patterns effectively, they often failed to generalize under long-range dependencies or concurrent branches. These challenges highlighted the need for architectures that process the entire trace as a structured object.
Building on this insight, researchers have developed global sequence architectures that process entire traces simultaneously. Transformer based models (bukhsh2021processtransformer; wuyts2024sutran; amiri2024pgtnet) leverage self attention mechanisms to capture long range dependencies, demonstrating improved performance over prefix based methods. Hierarchical variants (ni2023predictive) segment event sequences to capture multilevel semantics, addressing the complexity of nested or multigranular process structures. In parallel, global graph based approaches provide structured alternatives. For example, pasquadibisceglie2024prophet apply heterogeneous Graph Attention Networks (GATs) to model interactions across multiple edge types, while rama2023embedding combine Recurrent Neural Networks (RNNs) with GCNs to jointly capture temporal and structural dependencies. Although these models improve performance, they still lack clearly defined and explainable mechanisms for the dynamics among node positions, temporal relevance and semantic transitions.
@@ 2.2 Temporal Dynamics in Process Prediction
While many models treat transitions between events uniformly, recent graph based approaches have begun to emphasize their semantic diversity and predictive importance. pasquadibisceglie2024prophet incorporate multi type edges to distinguish between control flow and attribute based connections, demonstrating that transition semantics contribute significantly to prediction accuracy. ayaz2024semi apply graph autoencoders with edge conditioned convolutions to encode transition semantics, showing that different transition types require distinct representations. lischka2025directly preserve transition multiplicity using multi edge graphs and propose edge centric GNNs that treat transitions as first class entities. Despite these advances, the integration of semantic transition information into attention mechanisms remains rare and underexplored in current literature.
Recent work has also highlighted the role of temporal stability (teinemaa2018temporal) in shaping prediction accuracy, with studies showing that elapsed time between events carries complementary information beyond structural order (van2011time). Models that incorporate time aware embeddings (nguyen2020time) demonstrate clear benefits in handling heterogeneous case duration and non-uniform inter-event gaps (verenich2019survey). Yet, most existing approaches either treat time and transition semantics in isolation or apply them heuristically, leaving open the question of how to jointly integrate temporal relevance and semantic diversity into a unified attention framework.
@@ 2.3 Interpretability in PBPM
As predictive models in PBPM become more complex, interpretability has become essential for stakeholder trust and actionable insight. Early efforts relied on post hoc explainability methods such as surrogate models and feature attribution techniques (mehdiyev2021explainable; el2022xai), but these typically offered limited visibility into the actual decision making processes of deep neural models. Recent research has shifted toward exploiting intrinsic interpretability within graph neural networks and attention based architectures. The PROPHET framework (pasquadibisceglie2024prophet) integrates GNNExplainer to identify influential nodes and edges, while kappel2025explaining analyze attention weights in Transformer models to visualize control flow relevance across traces. Despite these developments, interpretability in GNN models for PBPM remains underdeveloped. Most existing work offers limited transparency into how temporal or structural signals shape predictions, and few approaches provide multilevel explanations that support both technical model validation and process level decision support.
@@ 3 Preliminaries and Definitions
Unlike prior PBPM graph formulations that focus only on event labels and direct-flow connectivity, our construction augments edges with both local and global temporal attributes as well as semantic embeddings. Local temporal gaps are represented as edge weights, while global temporal distances are encoded as a separate vector aligned with all edges. In addition, categorical transition types are embedded into higher-dimensional vectors, making edge attributes multi-dimensional rather than scalar. Together with the inclusion of sequence-level features $\mathbb{F}_{j}$, this design preserves global context while enriching the expressive power of the event-sequence graph.
@@ 4.1 Modeling Pipeline Overview
Our modeling framework evolves through a four stage pipeline, each stage progressively enhancing the models’ capacity to capture temporal and semantic dependencies in event sequences. Figure 1 provides a global overview of the unified architecture, highlighting how local–global comparison, temporal decay, and transition semantics are integrated across stages.
@@
Figure 1: Overview of the proposed framework. Event logs are transformed into attributed event-sequence graphs capturing local time gaps ($\Delta t$), global time decay ($\boldsymbol{\delta}$), and transition semantics ($\tau$). Prefix-based GCNs model local temporal dependencies across multiple prefix sizes, while full-trace GATs progressively incorporate temporal and semantic cues through layered attention mechanisms. The outputs correspond to next-event predictions for GCNs and sequence of next-event generation for GATs, with interpretability provided in the dynamic temporal (decay) based variants.
@@ 4.2.1 Graph Convolutional Networks
We adopt the GCNConv operator proposed by kipf2016semi, which uses a normalized adjacency matrix to stabilize message passing. The propagation rule at layer $l$ is adapted to our edge weighted, attributed graphs as follows:
@@ 4.3 Graph Attention Networks with Full Sequence Graphs
GATs enable flexible message passing by dynamically learning attention weights between connected nodes velickovic2017graph. Unlike GCNs, which aggregate neighbor features via fixed edge weights, GATs learn attention scores that reflect the relative importance of each neighbor during aggregation.
@@ 4.3.3 GAT Models with Full Graph
In addition to node attributes and local time gaps ($\Delta t$) as edge weights, each directed edge is also assigned a categorical transition-type identifier ($\tau$) that encodes the semantic relation between consecutive events. These identifiers are embedded into continuous vectors before entering the attention computation, enabling the model to learn how attention dynamics shift when transition semantics are introduced. The interaction between node position, temporal attributes, and transition embeddings is further analyzed and visualized in the case study (Appendix A) and through edge attention correlation analysis (Section 6.3.2).
@@ 4.3.4 Summary
Prior work has shown that temporal and semantic cues capture complementary aspects of sequential behavior. Local inter-event gaps ($\Delta t$) are widely used to encode short-range ordering (venkateswaran2021robust), while transition types ($\tau$) provide semantic context that improves process modeling and predictive accuracy (comuzzi2024process). Global temporal decay ($\delta$) has also been employed to model recency effects (wahid2019predictive), ensuring that more recent events carry greater influence. Our framework builds on these insights by systematically combining these cues within GAT variants: each variant highlights a distinct factor, and the final design (GAT-TDTE) unifies them in a principled, layered progression. Empirical results in Section 6 confirm that this design yields complementary gains rather than arbitrary stacking.
Unlike prefix-based GCNs, the GAT models operate on full event sequence graphs, capturing long-range dependencies across entire traces. Within this global setting, the variants provide a controlled comparison of how temporal gaps, decay mechanisms, and semantic transitions individually and jointly shape attention.
We first establish GAT-T as a direct baseline against prefix GCNs: both use local temporal gaps $\Delta t$, but GAT-T has access to the full trace, with $\Delta t$ capturing short-range ordering. GAT-TT extends this baseline by adding transition embeddings $\tau$, so that attention reflects not only temporal proximity but also the semantic type of each edge. This semantic encoding also enhances interpretability. Because each $\tau$ corresponds to a discrete process transition category (e.g., “approval to rejection” or “submission to revision”), the learned attention weights can be directly mapped to meaningful relational patterns within the process. This enables clear visualization of how edge-level semantics influence the temporal distribution of attention and reshape the model’s predictive focus (see Appendix A). Similar semantic-edge explainability strategies have been shown to improve transparency in process-aware GNNs (pasquadibisceglie2024prophet; lischka2025directly). To capture the recency effect observed in temporal data, we then move to dynamic temporal windows. GAT-TD replaces local $\Delta t$ with a decay term based on global distance $\delta$, producing recency-aware attention that downweights older events and creates case-specific attention windows. Finally, GAT-TDTE combines decay and transition semantics, integrating both *when* an event occurred (recency) and *what* transition it represents (semantics).
Interpretability signals per variant
Variant $\Delta t$ $\tau$ $\delta$ Attention logit (pre-softmax)
GAT-T $\checkmark$ $z(\cdot;\,\Delta t)$
GAT-$T$ $\checkmark$ $\checkmark$ $z(\cdot;\,\Delta t,\tau)$
GAT-TD $\checkmark$ $z(\cdot)\,e^{-\lambda\delta}$
GAT-TDTE $\checkmark$ $\checkmark$ $(z(\cdot)+s(\tau))\,e^{-\lambda\delta}$
Notes. $z(\cdot)$ is the LeakyReLU logit from $W\mathbf{v}_{i}\|W\mathbf{v}_{i+1}]$. Edges are $(i\!\to\!i{+}1)$.
Figure 2: Variants arranged by temporal cue (rows: local $\Delta t$ vs global decay $\delta$) and transition semantics (columns: without vs with $\tau$). Arrows denote incremental extensions: GAT-T $\xrightarrow{+\tau}$ GAT-$T$; GAT-T $\xrightarrow{\Delta t\to\delta}$ GAT-TD; GAT-TD $\xrightarrow{+\tau}$ GAT-TDTE (final).
Interpretability also grows layer by layer. Each added component introduces a distinct, inspectable term in the attention logit: raw logits $z$, local gaps $\Delta t$, global decay factors $e^{-\lambda\delta}$, and semantic embeddings $\tau$. Visualizing these terms (Fig. 2) shows how attention shifts from local timing to semantics, from local to global decay, and finally to their integration. This layered design makes the dependencies among variants explicit and demonstrates that the fusion of temporal and semantic signals is structured rather than simple feature stacking.
@@ 5.1 Data Description and Preprocessing
We evaluate our models on six benchmark event logs: BPI12 (vanDongen2012BPI), BPI12W (vanDongen2012BPI), BPI13i (Steeman2013BPI), BPI13c (Steeman2013BPIclosed), BPI20 (vanDongen2020BPI), and Helpdesk (Polato2017HelpDesk). These datasets originate from real-world business processes across multiple industries. BPI12 and BPI12W record loan applications from a Dutch financial institution, with BPI12W including work item logs. BPI13i (incident) and BPI13c (closed problem) stem from Volvo IT’s service management system. The Helpdesk log captures ticket handling in an Italian software company, while BPI20 (prepaid travel cost) covers two years of travel reimbursement requests from a Dutch university, including both domestic and international trips with multi-level approvals. Together, these datasets span diverse temporal scales, process complexities, and attribute richness (Table 1), providing a comprehensive testbed for our modeling pipeline.
@@
  • High Class Cardinality: BPI20 introduces a demanding setting with 29 prediction classes but only 2,099 traces. Its sequences are of moderate length (median 8, max 21), with minimal event-level features but several sequence-level attributes. This profile stress-tests model generalization when class space is large and per-event signal is sparse.
@@
Dataset #Seq Min(len) Max(len)33footnotemark: 3 Med(len) Min(t) Max(t) Med(t) #Attr(E, S) #Evt #St #Cls44footnotemark: 4
@@
BPI20 2099 1 21(18) 8 0 467965 34624 (1, 4) 29 1 29
11footnotetext: When only one sequence exists at the maximum length, we report the second-longest length in brackets (representing the maximum length found in the test dataset).22footnotetext: len = sequence length. t = time (minutes) between first and last events. #Attr(E, S) = number of event- and sequence-level attributes. med = median
A key observation from our dataset analysis is the discrepancy between the number of unique event types (#Evt) and the final classification labels (#Cls), which results from incorporating “lifecycle:transitions” as sub-status information (#St). For example, in the BPI13i and BPI13c datasets, the sub status categories (13 and 7, respectively) substantially increase the label space beyond the base event vocabulary (both with 4 activity types), leading to 3.25× and 1.75× larger output spaces. Given this expanded granularity in many real-world logs, we adopt the joint event and sub status prediction formulation, more commonly used in recent PBPM graph based models (rama2021deep; rama2023embedding). This design better reflects the multi-dimensional nature of process states and supports more fine grained, actionable predictions.